TEO Home
Home
Meetings
People
Publications

Pictures

Links

1Latest News

107/13/2009
RET Schedule of First Day ( Click here to download the schedule)

Attending:
Teacher participants: Donn Arnold, KISD, Dawn Chegwidden, LISD, Laura De Lemos, CFBISD, Sharon Wood, LISD. (Note: Laura will not be able to attend the first week and will start the program earlier)

Faculty Mentors: M. Acevedo, EE, R. Thompson, Biol., S. Fu, EE, Y. Huang, CSCE, X. Li, EE.

Dept. and College: M. Varanasi (EE Dept. Chair), M. Garcia-Rubio (Associate Dean)

Graduate students: J. Yang (EE and CSCE), S. Gurung (EE), C. Zhang (CSCE), N. Xu (EE and CSCE), K. Pathapati (EE and CSCE).

Undergraduate Students: Mitchell Horton (EE)

Schedule

10:00-10:30 AM Pre-assessment test. R. Thompson

10:30-10:45 AM Introductions and group dynamics. M. Acevedo and Y. Huang

10:45-11:00 AM Welcoming words on behalf of College of Engineering (CNEG) and overview of CENG outreach goals and activities M. Garcia-Rubio

11:00-11:15 AM Welcoming words on behalf of EE Dept. and overview of EE Dept. goals and activities O. Garcia

11:15-11:30 AM NSF's RET Concept and overall context of this RET in EE's NSF funded EE project "A Project-and Design-oriented Innovative Electrical Engineering Program" O. Garcia

11:30-Noon Orientation to Discovery Park. S. Fu

Noon-1:00 PM Lunch break. Sandwich lunch provided by EE.

1:00-1:15 PM Environmental context of this RET. CI-TEAM and CRI NSF-funded projects. M. Acevedo and Y. Huang

1:15-2:00 PM Presentations by faculty mentors and graduate students. Overview of their research.

2:00-2:45 PM Research methodology and work plan for this RET. J. Yang and X. Li

2:45-3:00 PM Group Discussion.

3:00-4:00 PM Tour of EE and CSCE facilities. Y. Huang, X. Li

102/18/2009
The NEON presentation of Jan that was cancelled due to inclement weather has been rescheduled for Thursday March 26 4:00 PM to 5:30 PM in EESAT 320A.. Please mark your calendars. It will be followed by a field visit to LBJ Friday morning March 27.

102/05/2009
Today we had a meeting with CAS Computer Service Team and Telcommunication Department in UNT. We will apply for 14 static IP addresses in a separate subnet for our project use. For system migration, we have finished migrating the web server, and are still in the process of testing new SBCs functioning as data collection server.

101/15/2009
Today we have finished syncronization process for database server TEO3. At this time TEO3 is running in parallel with TB3(old database server). We will evaluate its performance and reliability and then turn off TB3 in a weel also. Meanwhile the new data collection SBC is still under testing. New firewall modules need to be added. And since its JDK is not standard, we need further testing to confirm its compatability with existing programs.

110/30/2008
Dr. Acevedo has invited Prof Ermanno Pietrosemoli from ULA, Venezuela, to visit with us for one week Dec. 1-5. The purpose is to help us setup data link for TEO stations based on low cost Wifi and Wimax technology. He has a lot of practical experience doing this in the field. We can use CI-TEAM grant funds to pay for the airline ticket and per diem. Please go to http://wndw.net/ for info on his activities. They have a book explaining their methods.

109/18/2008
Dr. Fu created a web page at http://www.ee.unt.edu/public/fu/Project/cri_dev.php. From this link, we can see the information on all components we have tried, the item number, picture, price, and web link. As some of the information is sensitive, the page is password protected. Please check your email for password.

109/03/2008
We have a great trip this afternoon on the two sites: Pecan Creek Water Reclamation Plant and Hickory Creek Lift Station. We have been working on Greenbelt Park for more than 2 years. This is the first time we explore the new sites, which will be used for the new sensor networks.
Upon the information from this visit, the Pecan site maybe a good option for the deployment of large scale sensor networks. While the land is covered by some trees, most of space is open and has plenty of sun light, which is very important to our solar-driven sensor motes. Also, the AC power is available for base station. Comparing with Greenbelt, this site is very easy to access. We can drive there with our own car.
The Hickory Creek Site may be used for small scale sensor networks. It is also equipped with AC power. Some sensors are already deployed there, which may require some work to integrate them into our sensor system. The sensor motes (6-8) can be deployed along the fence near the creek. Check out latest pictures and videos at here.

108/31/2008
Today we visited Greenbelt station again to update SBC program, and started measuring sensor current. We also recorded the coordinates of the sensors, available here. This will be shown in a google map based interace at TEO intenal site. Check out latest pictures and videos at here.

108/30/2008
TEO SOS Service (for UV data) is available for test at here.

107/31/2008
Latest solar data collected by Adam is now available!

Get it here!

"Run 1" was under a plant and just for one day. "Run 2" is under a bush, where it should get an hour or two of direct sunlight in the afternoon, and the data seems to bear this out. There's a graph showing the current from the solar panel, in which you can clearly see the daily cycle. For this four-day period, the average current collected was 5.8mA.

106/23/2008
Jesse Heers, the technical sales representative of Campbel Scientific Inc., discussed with CRI team member on June 23. The CRI project plans to develop the environmental monitoring system with latest datalogger CR1000 for data acquisition. Attendees include Jerry, Miguel, Xinrong, and Shengli.

105/29/2008
The CRI group is working on the field test of wireless sensor networks. On May 29, Jerry, Adam, Sanjaya, and Dr. Fu visited Greenbelt park to deploy and test wireless sensor networks. The goal of this visit is to evaluate the effect of foliage and trees on the performance of solar panel.

Sensor Package
1.Sensor Package

Adam & Jerry
2. Adam & Jerry

Sanjaya
3. Sanjaya

102/13/2008
Greenbelt Weather Station SBC has been redeployed.

 

 

 

 

 

 


More News...

   Contact Us